please answer each question in its entirety, thanks!

1. When marginal product (MP or MPP) is falling, Average Product (AP or APP) and Total Product (TP) also must be falling. True or false? Explain.

2. Assuming output is increasing, if Total Fixed Costs (TFC) is constant, then Average Fixed Costs (AFC) must be falling. True or false? Explain.

3. If quantity of output is 100, TFC = $1,000, and Average Variable Cost (AVC) = $40, then Average Total Cost (ATC) must equal $140. True or false? Explain.
Answer #1

Ans 1)

Marginal product is additonal product produced for every additional unit of input.

If marginal contribution per unit of input is increasing this will increase rhe average product and decreaes when marginal product decreases for additional unit of input

Therefore statement is true.

Ans 2)

Average Fixed Cost =FC/Output

Fixed cost is constant but output is increasing therefore Average fixed cost will decrease

Hence statement is true

Ans 3)

Average total cost=Average fixed cost+Average variable cost

Average fixed cost=FC/Q=1000/100=10

Average variabke cost=40

Average total cost=40+10=50

Therefore statement is false.